# A Generative Modeling / Physics-Informed Neural Network Approach to Random Differential Equations

### v1

@@ 6.1 Well-Definedness of $U(\xi)$ Distribution
Based on the Random PDE stated in (6), we are interested in approximating its probability distribution, namely $\left[U\right]_{\!\#}\gamma$. We would like to know under what assumptions can we ensure that this distribution is well-defined. Moreover, does $\left[U\right]_{\!\#}\gamma$ depend on the choice of the underlying probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$?
@@
$G:\tilde{\mathcal{H}}\times\overline{\mathcal{H}}\times\Xi\to\mathcal{H}\text{with }\left(\tilde{f},\tilde{g},\tilde{\xi}\right)\mapsto\tilde{u}(\cdot;\tilde{\xi})\,,$
where $\tilde{u}(\cdot;\xi)$ is the solution to the Random PDE (similar to the one in (6)) for a fixed $\xi=\tilde{\xi}$, and with right hand side replaced by $\tilde{f}$ and $\tilde{g}$.
@@
  • The solution operator $G$ is well-defined for every point in ~ׯ×$\tilde{\mathcal{H}}\times\overline{\mathcal{H}}\times\Xi$, and
@@
Under these conditions, it follows that for any two probability spaces $\left(\Omega,\mathcal{F},\mathbb{P}\right)$ and $\left(\tilde{\Omega},\tilde{\mathcal{F}},\tilde{\mathbb{P}}\right)$, with corresponding random elements $\xi:\Omega\to\Xi$ and $\tilde{\xi}:\tilde{\Omega}\to\Xi$, both having law $\gamma$, we have:
$\left[U\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\right]_{\!\#}\gamma=\left[G\circ Q_{f,g}\circ\tilde{\xi}\right]_{\!\#}\tilde{\mathbb{P}}=\left[U\circ\tilde{\xi}\right]_{\!\#}\mathbb{P}\,.$
@@
6.2 Discrete Neural Network Spaces
@@
$v_{\theta}(\overline{y})=\mathcal{C}_{L}(\overline{y}):=C_{L}\circ\sigma\circ C_{L-1}\cdots\circ\sigma\circ C_{1}(\overline{y})\quad\forall\overline{y}\in\mathbb{R}^{d_{I}}.$ (29)
@@
$C_{k}y=W_{k}y+b_{k},\qquad\text{where }W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},b_{k}\in\mathbb{R}^{d_{k+1}},$ (30)
@@
The set of networks $\mathcal{C}_{L}$ with a given architecture of the form (29), (12) is called $\mathcal{N}.$ Then $\Theta\in\mathbb{R}^{\dim{\mathcal{N}}}$ where the total number of degrees of freedom of $\mathcal{N},$ is $\dim{\mathcal{N}}=\sum_{k=1}^{L}d_{k+1}(d_{k}+1).$ The nonlinear discrete set of functions $V_{\mathcal{N}}$ is defined as
$V_{\mathcal{N}}=\{u_{\theta}:\varOmega_{T}\to\mathbb{R},\ \text{where }u_{\theta}(\bm{x})=\mathcal{C}_{L}(\bm{x}),\ \text{for some }\mathcal{C}_{L}\in\mathcal{N}\,\}\,.$ (31)
6.3 Neural Networks with Random Input as Generative Models
In generative modeling, neural networks receive both a standard input and a random input (often called the latent code or noise). Formally, assume a probability distribution $\gamma$ over $\mathbb{R}^{d_{random}}$ (also known as the seed or reference distribution), and a neural network $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$. Then, the generative model is the random variable:
@@ 6.3 Neural Networks with Random Input as Generative Models
$\mu_{\theta}(x,\,A):=\gamma\circ Q_{\theta,x}^{-1}(A)=\int_{\mathbb{R}^{d_{\text{random}}}}\mathds{1}_{\{g_{\theta}(x,\xi)\in A\}}\,\gamma({\rm d}\xi)\,,\qquad\forall A\in\mathcal{B}\left(\mathbb{R}^{d_{output}}\right).$
If the mapping $g_{\theta}:\mathbb{R}^{d_{input}}\times\mathbb{R}^{d_{random}}\to\mathbb{R}^{d_{output}}$ is measurable with respect to the Borel $\sigma-$algebras $\mathcal{B}\left(\mathbb{R}^{d_{input}}\times\mathbb{R}^{d_{random}}\right)$ and $\mathcal{B}\left(\mathbb{R}^{d_{output}}\right)$, then the following properties hold:
  • For every fixed $x\in\mathbb{R}^{d_{input}}$, $\mu_{\theta}(x,\cdot)$ is a probability measure on $\mathbb{R}^{d_{output}}$
  • For every fixed $A\in\mathcal{B}\left(\mathbb{R}^{d_{output}}\right)$, the function $x\mapsto\mu_{\theta}(x,\,A)$ is measurable with respect to $\mathcal{B}\left(\mathbb{R}^{d_{input}}\right)$.
  • $\mu_{\theta}(\cdot,A)$ is $\mathcal{B}\left(\mathbb{R}^{d_{input}}\right)/\mathcal{B}\left(\mathbb{R}^{d_{output}}\right)-$measurable
Therefore, $g_{\theta}$ defines a stochastic kernel $\mu_{\theta}(x,\,{\rm d}y)$ from $\mathbb{R}^{d_{input}}$ to $\mathbb{R}^{d_{output}}$.
@@
$\mathcal{B}\left(\mathbb{R}^{d_{input}}\times\mathbb{R}^{d_{random}}\right)=\mathcal{B}\left(\mathbb{R}^{d_{input}}\right)\otimes\mathcal{B}\left(\mathbb{R}^{d_{random}}\right)\,,$
since both $\mathbb{R}^{d_{input}}$ and $\mathbb{R}^{d_{random}}$ are Polish spaces, see Theorem 17.25 in [kechris2012classical].
[…]

### v2

@@ A.1 Well-Definedness of $U(\xi)$ Distribution
Based on the Random PDE stated in Equation 6, we are interested in approximating its probability distribution, namely $\left[U\right]_{\!\#}\gamma$. We would like to know under what assumptions can we ensure that this distribution is well-defined. Moreover, does $\left[U\right]_{\!\#}\gamma$ depend on the choice of the underlying probability space $\left(\Omega,\mathcal{F},\mathbb{P}\right)$?
@@
[…]
where $\tilde{u}(\cdot;\xi)$ is the solution to the Random PDE (similar to the one in Equation 6) for a fixed $\xi=\tilde{\xi}$, and with right hand side replaced by $\tilde{f}$ and $\tilde{g}$.
@@
  • The solution operator $G$ is well-defined for every point in $\tilde{\mathcal{H}}\times\overline{\mathcal{H}}\times\Xi$, and
@@
[…]
$\left[U\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\circ\xi\right]_{\!\#}\mathbb{P}=\left[G\circ Q_{f,g}\right]_{\!\#}\gamma=\left[G\circ Q_{f,g}\circ\tilde{\xi}\right]_{\!\#}\tilde{\mathbb{P}}=\left[U\circ\tilde{\xi}\right$$_{\!\#}\mathbb{P}$.
@@
A.2 Discrete Neural Network Spaces
@@
(43) $v_{\theta}(\overline{y})=\mathcal{C}_{L}(\overline{y}):=C_{L}\circ\sigma\circ C_{L-1}\cdots\circ\sigma\circ C_{1}(\overline{y})\quad\forall\overline{y}\in\mathbb{R}^{d_{I}}.$
@@
(44) $C_{k}y=W_{k}y+b_{k},\qquad\text{where }W_{k}\in\mathbb{R}^{d_{k+1}\times d_{k}},b_{k}\in\mathbb{R}^{d_{k+1}},$
@@
The set of networks $\mathcal{C}_{L}$ with a given architecture of the form Eq. 43, Eq. 12 is called $\mathcal{N}$. Then $\Theta\in\mathbb{R}^{\dim{\mathcal{N}}}$ where the total number of degrees of freedom of $\mathcal{N},$ is $\dim{\mathcal{N}}=\sum_{k=1}^{L}d_{k+1}(d_{k}+1).$ The nonlinear discrete set of functions $V_{\mathcal{N}}$ is defined as
(45) $V_{\mathcal{N}}=\{u_{\theta}:\varOmega_{T}\to\mathbb{R},\ \text{where }u_{\theta}(\bm{x})=\mathcal{C}_{L}(\bm{x}),\ \text{for some }\mathcal{C}_{L}\in\mathcal{N}\,\}.$
A.3 Neural Networks with Random Input as Generative Models
In generative modeling, neural networks receive both a standard input and a random input (often called the latent code or noise). Formally, assume a probability distribution $\gamma$ over $\mathbb{R}^{d_{\text{ran}}}$ (also known as the seed or reference distribution), and a neural network $g_{\theta}\in\mathcal{V}_{\mathcal{N}}$. Then, the generative model is the random variable:
@@ A.3 Neural Networks with Random Input as Generative Models
$\mu_{\theta}(x,\,A):=\gamma\circ Q_{\theta,x}^{-1}(A)=\int_{\mathbb{R}^{d_{\text{ran}}}}\mathds{1}_{\{g_{\theta}(x,\xi)\in A\}}\,\gamma({\rm d}\xi),\qquad\forall A\in\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right).$
If the mapping $g_{\theta}:\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\to\mathbb{R}^{d_{\text{out}}}$ is measurable with respect to the Borel $\sigma$-algebras $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\right)$ and $\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$, then the following properties hold:
  • For every fixed $x\in\mathbb{R}^{d_{\text{in}}}$, $\mu_{\theta}(x,\cdot)$ is a probability measure on $\mathbb{R}^{d_{\text{out}}}$
  • For every fixed $A\in\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$, the function $x\mapsto\mu_{\theta}(x,\,A)$ is measurable with respect to $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)$.
  • $\mu_{\theta}(\cdot,A)$ is $\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)/\mathcal{B}\left(\mathbb{R}^{d_{\text{out}}}\right)$-measurable
Therefore, $g_{\theta}$ defines a stochastic kernel $\mu_{\theta}(x,\,{\rm d}y)$ from $\mathbb{R}^{d_{\text{in}}}$ to $\mathbb{R}^{d_{\text{out}}}$.
@@
$\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\times\mathbb{R}^{d_{\text{ran}}}\right)=\mathcal{B}\left(\mathbb{R}^{d_{\text{in}}}\right)\otimes\mathcal{B}\left(\mathbb{R}^{d_{\text{ran}}}\right),$
since both $\mathbb{R}^{d_{\text{in}}}$ and $\mathbb{R}^{d_{\text{ran}}}$ are Polish spaces, see Theorem 17.25 in [20].
Figure 4: Training and testing loss over time during the training of the PINN architecture for the bistable ODE Eq. 19.
Figure 5: Training and testing loss function over time of the training of the PINN and the PINN-PCE architecture for the diffusion equation Eq. 21.
Figure 6: Training and testing loss function over time of the training of the PINN and the PINN-PCE architecture for the reaction-diffusion equation Eq. 27.
Acknowledgments
The authors used AI-assisted tools (ChatGPT by OpenAI) for minor editorial assistance, including grammar correction and language polishing. All substantive ideas, analysis, and results presented in this work are entirely the authors’ own.